4 consecutive odd numbers have an average of 25. What are the 4 integers?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

Let x, x+2, x+4, x+6e the 4 numbers.
(x+x+2+x+4+x+6)/4=25
(4x+12)/4=25
4x+12=100
4x=100-12
4x=88
x=88/4
x=22 ans. for the smallest number.
22+2=24 for the next number
22+4=26 for the next number.
22+6=28 for the largest number.
Proof:
(22+24+26+28)=/4=25
100/4=25
25=25

Answer by Alan3354(69443)

There are no odd numbers that fit.